7x-8y=9
solve for y

Answer by alex224(8)


To solve for y you need to get the y isolated.
The first step is to move the x over to the opposite side of the equation by using subtraction.
-8y=9-7x

Then to get the y alone you need to devide both sides by -8
-8y/-8=9-7x/-8

The final answer is going to be
y=-9+7x/8